OpenStudy (anonymous):

Another partial fraction question, this time they gave me int 1/(x^2+xsqrt(x))dx, how do I do it when I only have a constant on top but nothing that would give me a constant on the bottom?

OpenStudy (anonymous):

In equation form: \[\int\limits_{}^{}\frac{ 1 }{ x^2+x \sqrt{x} }dx\]

hartnn (hartnn):

lets simplify this by first substituting sqrt x as y so whats x^2 =... ? x\sqrt x = ... ?

OpenStudy (anonymous):

I'm sorry, I'm still lost by that

hartnn (hartnn):

\(\sqrt x=y\) square both sides, what u get ?

OpenStudy (anonymous):

x=y^2

hartnn (hartnn):

right so, x^2 = ... ?

OpenStudy (anonymous):

y^4, but im not seeing how this is helping yet because there's still a dx in the equation

hartnn (hartnn):

right, that will help only for partial fractions part! so we have y^4 + y^3 in the denominator right ?? which is y^3 (1+y) can u do partial fractions for this ?

hartnn (hartnn):

in the end, just replace back, y =sqrt x

OpenStudy (anonymous):

yeah but if i did the substitution like you're suggesting now i'd have \[\int\limits_{}^{}\frac{ 1 }{ y^4+y^3 }dx\] and you can't mix variables so what happens to the dx?

OpenStudy (anonymous):

oh wait you mean just use the y to simplify the algebra and then sub back in with the dx, i get it now

hartnn (hartnn):

as i just said, use it for partial only forget about the integral and dx for now

hartnn (hartnn):

yes yes!